OpenStudy (anonymous):

Will Medal Solve the system of equations. x + 3y = −1 2x + 2y = 6 (−4, 1) (2, −1) (4, −1) (5, −2)

OpenStudy (anonymous):

@boldjon

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (neonumbrella5115):

Do you have a guess? Or have you tried to solve it?

OpenStudy (anonymous):

yeah my guess was b

OpenStudy (trojanpoem):

-2x - 6 y= 2 2x + 2y = 6 -4 y = 8 y = -2 x = - 1 - 3y = -1 - 3 * -2 x = -1 + 6= 5 (5,-2)

OpenStudy (anonymous):

so its d then

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

True.

OpenStudy (anonymous):

k thx